How many grams of silver nitrate would you need to dissolve in 500 of water in order to make a 6.0 M solution ? Show your work .


1
Expert's answer
2021-02-01T03:57:10-0500

number of moles = volume*concentration.

n = (500 ml)*(6 mol/L) = (0.5 ml)*(6 mol/L) = 3 mol

m = n*Mr = (3 mol)*(169.87 g/mol) = 509.61 g